At what point do the curves r1(t) = <t,1-t,t> and r2(t)=<3+s, s-2, (s^2)+3> intersect?

Find their angle of intersection
Answer #1

Given, 7,16) (t,i-t, t> 72 (8) = {3+S, 5-2, 5²+3) at intersection point, t=375 ; i-tes-d; t=5+3 = ) 5+3 = 375 t = 3+S ; t = 3direction of y, 1t) = x/lt) - <1-11) X C3) = {1,-1,0) directin of yz(s) = x;(s) = (1,1,28) 2.(0) = <!,1,07 ii Coso- 8, (3)